Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
garaventa, sorrentino - 17:24 Thursday 02 December 2021 (54078) Print this report
SC AA loop: about 99% of MM reached

This morning shift we worked on the SC AA loop.

  1. First of all we have replaced the cable of RF_PD in order to reduce power losses (photo attached).
  2. We checked Power on SC beam path:

    19 mW out of fiber 

    18 mW after power control PBS

    17.5 mW before FI

    15 mW first pass after FI (towards OPA)

    14 mW second pass on FI (towads HD)

    12 mW reflected from IR shutter towards AA PD

    0.8 mW after filter before AA PD

  3. In this condition we had a filter in front of RF_PD and we investigated on the SC AA loop with DL on. The strategy was the following:
  • we moved the SC_PD_MIR in order to center beam on RF_PD with SC AA loop engaged (Err 4MHz I); then, AA loop opened,  we opened IR shutter and we looked at PD_MONI scanning OPA and Pump Phase. As you can see in Fig.2 we reached about 97 % of MM. We noticed that M3_X and M3_Y were saturating, so we discharged them and before we adjusted demodulation phase of M3_Y. 
  • After Actuators discharging and realigning by hand, we engaged SC AA loop improving the signal on RF_PD, and then we measured 99% of MM on PD_MONI, as you can see in Fig.3,4. We added the new setpoints of M1 and M3 on the configuration file EQB1_HD_AA.cfg (lines 141-144).

Images attached to this report
Comments to this report:
garaventa - 8:25 Friday 03 December 2021 (54088) Print this report

In attached plots with the SC AA signals.

Images attached to this comment
Search Help